Mwema plays a game in which a fair die is thrown once. If the score is 1,2or 3 Mwema loses £10. If the score is 4 or 5 Mwema wins £x..If the score is 6, Mwema wins £2x.


1
Expert's answer
2021-04-29T17:13:04-0400

The probability distribution for Mwema is given by-




Then Expected amount of wins/loss ="\dfrac{-10}{2}+\dfrac{x}{3}+\dfrac{2x}{6}"


"=\dfrac{2x}{3}-5"


Since He does not get any profit so,

"\dfrac{2x}{3}-5=0"


"\Rightarrow x=7.5"


Hence The value of x is 7.5.